# Learning Frequency-Aware Dynamic Transformers for All-In-One Image Restoration

### v1

@@ III-B Degradation-adaptive restoration transformer
where $\boldsymbol{z}^{{}^{\prime}}$ represent the rescaled attention map, and $\boldsymbol{M}_{k}$ denotes the scaling coefficient for the $k$-th frequency band. The direct component of $z$ serves as a baseline, remaining unscaled to provide a reference for scaling other frequency bands. The set of scaling coefficients, $\boldsymbol{M}=\{\boldsymbol{M}_{1},\ldots,\boldsymbol{M}_{L-1}\}$, is learned through a degradation projection process implemented by a two-layer MLP. This MLP takes the degradation representations $d$, derived from Dformer, as input. The MLP is initialized such that the values of $M$ are zero, resulting in $\boldsymbol{z}^{\prime}=\boldsymbol{z}$. Essentially, the attention map $z$ is decomposed into multiple frequency bands, each of which is scaled by a coefficient learned through a degradation-aware projection. This allows for adaptive restoration based on the degradation characteristics.
III-C Composite training loss
The training of our approach is carried out in two distinct stages. Initially, Dformer is trained to learn degradation representations with a contrastive learning loss $\mathcal{L}_{cl}$. We consider $\boldsymbol{d}$ as the degradation representation of the anchor sample, $\boldsymbol{d}^{+}$ and $\boldsymbol{d}^{-}$ as the degradation representation of positive and negative samples obtained through the MoCo framework, where positive samples and the anchor sample come from the same degraded image, while negative samples come from other degraded images. $\mathcal{L}_{cl}$ is defined by:
@@ III-C Composite training loss
where $\mathcal{L}_{rec}=\frac{1}{T}\sum_{i=1}^{T}|\hat{I}_{i})-y_{i}|$ is a L1 loss. Here, $\hat{I}_{i}$ denotes the recovered image through Rformer, and $y$ is the corresponding clean image.
@@ IV Experiments and Results
In this section, we comprehensively evaluate and analyze our methods across four tasks: denoising, deraining, dehazing, and deblurring.
@@ IV-A Experimental Setup
Datasets. Following the existing works [15, 18], we assess the effectiveness of the proposed approaches in multi-degradation restoration using seven datasets: BSD400 [41], BSD68 [41], WED [42], and Urban100 [43] for image denoising, Rain100L [44] for image deraining, RESIDE [45] for image dehazing, and GoPro [46] for image deblurring.
Implementation details. The training settings are outlined following AirNet [15]. AdamW is used as the optimizer. The training phase consists of $1000$ epochs: the first $100$ epochs train the Encoder with Contrastive Loss optimization for warm-up, and the remaining $900$ epochs optimize the entire network using total loss optimization. The learning rate starts at $43e-$ 4, reducing to $53e-$ 5 after $60$ epochs, and then starts at $41e-$ 4 for the remaining epochs, halving every $125$ epochs. We fix the image patch size at $128\times 128$ and apply random data augmentations. The batch size is set to $400\times N$, where $N$ is the number of degradation types. We set the number of frequency bands to $L=2$ to balance efficiency and performance, as demonstrated in Section VII.
@@ IV-B Comparison to the state-of-the-art
We first perform a comparison to the state-of-the-art in the conventional ”noise-rain-haze” setting to showcase the superiority of our approach. Our comparison encompasses four single-degradation image restoration techniques, namely BDRNet [47], LP-Net [48], FDGAN [49], and MPRNet [50], alongside the multi-task method for multiple degradation image restoration, DL [51]. We are also specifically evaluating two specialized all-in-one methods, AirNet [15] and PromptIR [18].
The results, as shown in Table I, highlight the superiority of all all-in-one methods over other baselines for single degradation, underscoring their ability to address various unknown degradations within a unified framework. Notably, our approach demonstrates even better performance compared to other all-in-one methods. Specifically, we surpass AirNet [15] across all tasks, achieving an average performance improvement of $1.12$ dB PSNR and $0.011$ SSIM. Furthermore, we outperform PromptIR [18] in denoising and deraining tasks, with an average performance improvement of $0.26$ dB PSNR and $0.008$ SSIM.
Qualitative examples are presented in Fig. 2. Compared to AirNet [15] and PromptIR [18], our approach better preserves edge details when performing denoising and deraining, and achieves better color fidelity during dehazing.
TABLE I: Comparison to the state-of-the-art on the conventional “noise-rain-haze” setting. Existing all-in-one methods surpass other baselines designed for single degradation tasks, whereas our approach achieves superior performance.
@@
(a) Degraded images
(b) AirNet
(c) PromptIR
(d) Ours
(e) Groundtruth
Figure 2: The performance of various methods on denoising of $\sigma=25$ (first row), deraining (second row), and dehazing tasks (last row). In the blue-highlighted regions, our method demonstrates superior edge detail preservation for both deraining and denoising tasks. Additionally, it achieves better color fidelity in the dehazing task compared to other methods.
@@ IV-C Comparison on the number of degradation types
In this experiment, we conduct a comparative analysis between the proposed method, AirNet [15], and PromptIR [18] across different numbers of degradations to assess the stability of our approach. The experimental results in Table II show that as the number of degradation types increases, the network’s ability to restore images diminishes, resulting in a performance decline. Notably, both AirNet and PromptIR experience clear performance degradation when tasked with handling multiple degradations simultaneously. For instance, the PSNR for deraining drops from 38.31 dB to 34.7 dB for AirNet, and from 39.32 dB to 36.14 dB for PromptIR, as the number of combined degradation types increases from 2 to 4. This decline in performance occurs due to potential conflicts between different tasks during joint learning, which AirNet and PromptIR struggle to manage effectively. In contrast, our method explicitly addresses task disparities in frequency domain through frequency-aware dynamic transformers. Consequently, our method experiences a milder drop of only 1.43 dB, from 39.51 dB PSNR to 37.35 dB PSNR, showcasing superior stability across varying numbers of degradation types.
TABLE II: Comparison on the number of degradation types. As the number of combined degradation types increases, our proposed approach demonstrates superior performance stability compared to AirNet and PromptIR.
@@ IV-D Results on various combined degradations
In this section, we examine the impact of various combinations of degradation types on model performance, as detailed in Table III. When we randomly select and combine two out of the four degradation types, we observe that denoising performance remains relatively stable, regardless of whether it is combined with deraining, dehazing, or deblurring. This stability is likely because the denoising task dominates the training process, benefiting from a larger dataset across three noise levels: $\sigma=15$, $\sigma=25$, and $\sigma=50$.
For the deraining task, performance is optimal when combined with denoising, compared to combinations with dehazing or deblurring. This is likely due to both deraining and denoising focusing on recovering high-frequency details, thus aligning their frequency optimization directions. Conversely, dehazing and deblurring aim to remove low-frequency content, and their performance is enhanced when combined with denoising, due to the larger training dataset. Similar trends are observed when we randomly select and combine three out of the four degradation types, further supporting these findings.
TABLE III: Results on various combined degradations. Tasks can enhance each other when their degradation types (e.g., deraining and denoising) share similar frequency optimization directions. Conversely, when degradation tasks (e.g., deraining and dehazing) have conflicting optimization goals, a performance drop is observed.
@@
IV-E Ablation Studies
In this section, we present the ablation experiments outlined in Table IV-E and IV-E to validate the effectiveness of the proposed Dformer and Rformer models, along with their individual components. These experiments are conducted under the standard ”noise-rain-haze” setting. For clarity and conciseness, we report only the average PSNR and SSIM metrics.
Dformer. The key components of Dformer are the Input Frequency Decomposition (IFD) and Frequency-Aware Transformer Blocks (FA-TB). For comparison, we used the Swinformer as a baseline, which incorporates standard Swin Transformer blocks. Additionally, we created a second baseline by combining our Input Frequency Decomposition (IFD) with Swinformer. Both baselines, along with our Dformer, utilize the Rformer for restoration. As shown in Table IV-E, incorporating IFD into the Swinformer results in a slight performance improvement, with the average PSNR increasing from 31.53 to 31.63. However, when replacing the standard Swin Transformer blocks with our FA-TB, the average PSNR further improves from 31.63 to 32.32. These results underscore the importance of IFD and FA-TB in learning better degradation representations and enhancing overall image restoration performance.
Rformer. Rformer is designed following the architecture of Uformer, so we use Uformer as the primary baseline. Uformer addresses all tasks simultaneously without leveraging any degradation priors. The key component of Rformer is the Degradation-Adaptive Self-Attention (DA-SA). DA-SA dynamically rescales different frequency bands of the attention map to achieve adaptive restoration, guided by degradation representations acquired from Dformer. To further evaluate DA-SA, we developed an additional baseline where the rescaling is performed using learnable parameters without any degradation guidance. As shown in Table IV-E, our Rformer achieves the best performance, demonstrating the importance of DA-SA in enhancing restoration by effectively adapting to degradation characteristics.
TABLE IV: Ablation study
of Dformer.
Methods Average Swinformer 31.53/0.914 Swinformer+IFD 31.62/0.915 Dformer 32.32/0.921
TABLE V: Ablation study
of Rformer.
Methods Average Uformer 31.01/0.902 Scaling Uformer 30.91/0.898 Rformer 32.32/0.921
IV-F Further analysis

### v2

@@ III-C Degradation-adaptive restoration transformer
where $\boldsymbol{z}^{{}^{\prime}}$ denotes the modulated attention map, $\boldsymbol{M}_{k}$ denotes the modulation ratio for the $k$-th frequency band. $\boldsymbol{M}_{k}$ is initialized to zero, ensuring that $\boldsymbol{z}^{\prime}=\boldsymbol{z}$, indicating that no modulation is applied initially. The direct component serves as a baseline for modulating other frequency bands since it remains unscaled. Put plainly, the modulation process entails segmenting the attention map into distinct frequency bands and independently modulating each band using specific modulation ratios. The frequency modulation ratios $M=\{\boldsymbol{M}_{0},\boldsymbol{M}_{1},\ldots,\boldsymbol{M}_{L}\}$ are embedded using a degradation projection implemented by a two-layer MLP, which takes the degradation representations $\boldsymbol{d}$, derived from Dformer, as input.
III-D Composite training loss
The training of our approach is carried out in two distinct stages. Initially, Dformer is trained to learn degradation representations with a contrastive learning loss $\mathcal{L}_{cl}$. We consider $\boldsymbol{d}$ as the degradation representation of the anchor sample, $\boldsymbol{d}^{+}$ and $\boldsymbol{d}^{-}$ as the degradation representation of positive and negative samples obtained through the MoCo framework, where positive samples and the anchor sample come from the same degraded image, while negative samples comes from other degraded images. The samples are random patches of size $128\times 128$ from the degraded images and undergo random data augmentation. $\mathcal{L}_{cl}$ is defined by:
@@ III-D Composite training loss
where $\mathcal{L}_{rec}=|\hat{\boldsymbol{I}}-\boldsymbol{Y}|$ is a L1 loss. Here, $\hat{\boldsymbol{I}}$ denotes the recovered image through Rformer, and $\boldsymbol{Y}$ is the corresponding clean image.
@@ IV Experiments and Results
In this section, we comprehensively evaluate and analyze our proposed method across five tasks: denoising, deraining, dehazing, deblurring, and low-light enhancement. We start by detailing the experimental setup, followed by presenting both qualitative and quantitative results from benchmark tests, which demonstrate the effectiveness of our method for all-in-one image restoration. Additionally, we conduct ablation studies to validate the contribution of each component of our approach. Finally, we provide further analysis to highlight additional benefits of our method.
@@ IV-A Experimental Setup
Datasets. To assess the effectiveness of the proposed approaches in multi-degradation restoration, we mainly employ seven datasets: BSD400 [68], BSD68 [68], WED [69], and Urban100 [70] for image denoising, Rain100L [71] for image deraining, RESIDE [72] for image dehazing, GoPro [73] for image deblurring, and LOL dataset [74] for low-light enhancement. Detailed information on the datasets is provided in Table I. Following Li et al. [15], the noisy images for image denoising are generated by manually adding white Gaussian noise to the clean images at three corruption levels, namely, $\sigma=15,25,50$.
TABLE I: The details of datasets used for training and testing. The number in each bracket indicates the quantity of images or image pairs contained in the corresponding dataset.
Degradation type Training Testing
Denoise BSD400 [68] (400), WED [69] (4744) BSD68 [68] (68)
Derain Rain100L-train [71] (200) Rain100L-test [71] (100)
Dehaze RESIDE-OTS [72] (72135) RESIDE-SOTS [72] (500)
Deblur GoPro-train [73] (2103) GoPro-test [73] ($1$)
Low-light LOL-train [74] (485) LOL-test [74] (15)
Implementation details. The Rformer in this paper represents a general image restoration architecture based on the Transformer framework, adaptable to various specific Transformer implementations. In experiments, We leverage Uformer [43] as the Rformer, due to its extensively validated performance and computational efficiency across various restoration tasks. Our approach is implemented using the PyTorch framework and trained on a single NVIDIA A800 GPU.
During training, we utilize the AdamW optimizer to optimize the network parameters, following the configuration settings outlined in AirNet [15]. The training phase takes 1000 epochs: the initial 100 epochs are dedicated to training the Encoder using Contrastive Loss optimization for warm-up, while the subsequent 900 epochs focus on optimizing the entire network (both Encoder and backbone) using total loss optimization. In the first 100 epochs, the learning rate is initialized at 3e-4 and is then reduced to 3e-5 after the 60th epoch. For the subsequent 900 epochs, the learning rate is initialized at 1e-4 and is halved every 125 epochs to facilitate convergence. During training, we fix the image patch size at $128\times 128$ and apply random data augmentations to the input image pairs, in line with the standard practices of contrastive learning methodologies. The batch size is set to $400\times N$, where $N$ represents the number of degradation types. We set the number of frequency bands to $L=2$ across all experiments to achieve an optimal balance between efficiency and performance, as demonstrated by the experiments in Section X.
During testing, to ensure consistency in patch size for our Transformer-like architecture, we segment images into multiple $128\times 128$ patches, following the same procedure as in training. For images larger than $640\times 360$, we divide them into non-overlapping patches. For images smaller than or equal to $640\times 360$, we adopt the strategy from SwinIR, sampling a patch every $64\times 64$ pixels. This approach effectively balances border artifacts and computational efficiency.
@@ IV-B Comparison to the state-of-the-art
We first perform a comparison to the state-of-the-art in the conventional “noise-rain-haze” setting to showcase the superiority of our approach. Our comparison encompasses four single-degradation image restoration techniques, namely BDRNet [75], LP-Net [76], FDGAN [28], and MPRNet [52], alongside the multi-task method for multiple degradation image restoration, DL [77]. We are also specifically evaluating two specialized all-in-one methods, AirNet [15] and PromptIR [17].
The results, as shown in Table II, highlight the superiority of all all-in-one methods over other baselines for single degradation, underscoring their ability to address various unknown degradations within a unified framework. Notably, our approach demonstrates even better performance compared to other all-in-one methods. Specifically, we surpass AirNet [15] across all tasks, achieving an average performance improvement of $1.12$ dB PSNR and $0.011$ SSIM. Furthermore, we outperform PromptIR [17] in denoising and deraining tasks, with an average performance improvement of $0.26$ dB PSNR and $0.008$ SSIM.
Qualitative examples are presented in Fig. 4,5, and 6. Compared to AirNet [15] and PromptIR [17], our approach better preserves edge details when performing denoising and deraining, and achieves better color fidelity when performing dehazing.
(a) Degraded
(b) AirNet
(c) PromptIR
(d) Ours
(e) Clear
Figure 4: The performance of various methods on denoising ($\sigma=25$) tasks. From the regions masked by the blue rectangles, we observe our method better preserves edge details (best viewed digitally).
(a) Degraded
(b) AirNet
(c) PromptIR
(d) Ours
(e) Clear
Figure 5: The performance of various methods on deraining tasks. From the regions masked by the blue rectangles, we observe our method performs well, especially when recovering high-frequency details (best viewed digitally).
(a) Degraded
(b) AirNet
(c) PromptIR
(d) Ours
(e) Clear
Figure 6: The performance of various methods on dehazing tasks. From the regions masked by the green rectangles, we observe our method achieves better color fidelity (best viewed digitally).
TABLE II: Comparison to the state-of-the-art on the conventional “noise-rain-haze” setting. Existing all-in-one methods surpass other baselines designed for single degradation tasks, whereas our approach showcases superior performance.
@@ IV-C Comparison on the number of degradation types
In this experiment, we conduct a comparative analysis between the proposed method, AirNet [15], and PromptIR [17] across different numbers of degradations to assess the stability of our approach. The experimental results in Table III shows that as the number of degradation types increases, the network’s ability to restore images diminishes, resulting in a performance decline. Notably, both AirNet and PromptIR experience significant performance degradation when tasked with handling multiple degradations simultaneously. For instance, the PSNR for deraining drops from 38.31 dB to 34.7 dB for AirNet, and from 39.32 dB to 36.14 dB for PromptIR, as the number of combined degradation types increases from 2 to 4. This decline in performance occurs due to potential conflicts between different tasks during joint learning, which AirNet and PromptIR struggle to manage effectively.
In contrast, our method explicitly addresses task disparities in frequency domain through frequency-aware dynamic transformers. Consequently, our method experiences a milder drop of only 1.43 dB, from 39.51 dB PSNR to 37.35 dB PSNR, showcasing superior stability across varying numbers of degradation types. When four types of combined degradations are present, qualitative examples from deblurring tasks are shown in Fig. 7. Our approach demonstrates superior edge detail preservation compared to AirNet [15] and PromptIR [17].
TABLE III: Comparison on the number of degradation types. As the number of combined degradation types increases, our proposed approach demonstrates superior performance stability compared to AirNet and PromptIR.
@@
(a) Degraded
(b) AirNet
(c) PromptIR
(d) Ours
(e) Clear
Figure 7: The performance of various methods on deblurring tasks. From the regions masked by the blue rectangles, we observe our method better recovers high-frequency edges (best viewed digitally).
@@ IV-D Results on various combined degradations
In this section, we examine the impact of various combinations of degradation types on model performance, as detailed in Table IV. When we randomly select and combine two out of the four degradation types, we observe that denoising performance remains relatively stable, regardless of whether it is combined with deraining, dehazing, or deblurring. This stability is likely because the denoising task dominates the training process, benefiting from a larger dataset across three noise levels: $\sigma=15$, $\sigma=25$, and $\sigma=50$.
For the deraining task, performance is optimal when combined with denoising, compared to combinations with dehazing or deblurring. This is likely due to both deraining and denoising focusing on recovering high-frequency details, thus aligning their frequency optimization directions. Conversely, dehazing and deblurring aim to remove low-frequency content, and their performance is enhanced when combined with denoising, due to the larger training dataset.
Similar trends are observed when we randomly select and combine three out of the four degradation types, further supporting these findings.
TABLE IV: Results on various combined degradations. Tasks can enhance each other when their degradation types (e.g., deraining and denoising) share similar frequency optimization directions. Conversely, when degradation tasks (e.g., deraining and dehazing) have conflicting optimization goals, a performance drop is observed.
@@
(a) Denoising at different $\sigma$ values
(b) 4 D-types Image Restoration
Figure 8: Average modulation ratio on different degradation type. Learned, directly learned average modulation ratios under One-by-one setting. Embedded, average modulation ratios embedded from Dformer under All-in-one setting.
(a) PromptIR
(b) Uformer
(c) Ours
Figure 9: t-SNE visualization. Without frequency modulation, Uformer [43] fails to provide a discriminative capability superior to PromptIR[17]. However, with the incorporation of frequency modulation, our method effectively distinguish between different degradation types.
IV-E Analysis on Frequency Modulation
In this section, we analyze whether the frequency modulation of the attention map captures the aforementioned frequency characteristics of various degradations.
First, we let the modulation ratios in Rformer be learnable, instead of embedding them from Dformer. Under the One-by-one setting with L=2, we train 4 models for the denoising task under Gaussian noise levels of $\sigma$=15,25,50,75, respectively, and visualize the average of the learned modulation ratios, obtaining the ”Learned” entry in Fig. 8(a). Since we do not modulate the lowest frequency in our method, i.e., the DC component, the average of the modulation parameters here is the average modulation ratio for high frequencies. From this, we find that even without using our Dformer to capture frequency characteristics, the model can still adapt to the noise level requirements by adjusting the proportion of different frequency bands in the attention map. For all four noise levels, the average high-frequency modulation ratio is less than 0, showing a trend of suppressing high frequencies. As the noise level increases, the high-frequency noise becomes larger, so the model needs to suppress high frequencies more, and the modulation ratio becomes smaller. Of course, although directly learnable modulation ratios can capture the frequency differences of different degradation levels, it is not suitable for the all-in-one setting because it requires pre-specifying the degradation type to obtain the corresponding modulation ratios.
Then, we replace the learnable modulation ratios with those embedded by Dformer and performe the same experiment under the all-in-one setting. Since the modulation ratios are no longer model parameters but are embedded from the input degraded image, we sample 100 images for each of the four noise levels and average the resulting average modulation ratios, obtaining the ”Embedded” entry in Fig. 8(a), finding that it still shows the same trend. Finally, we performe the same experiment on the 4 tasks All-in-one of denoising, deraining, dehazing, and deblurring, obtaining the ”Embedded” entry in Fig. 8(b), finding that denoising and deraining aim to suppress high-frequency noise, as the average modulation ratios is less than 0, while dehazing and deblurring are the opposite. In Fig. 1 and Fig. 8, we can further find that even when comparing the dehazing and deblurring, which both restore high-frequency details, there are still some differences in frequency, that is, deblurring needs to restore high-frequency details more than dehazing.
To demonstrate that the proposed method not only captures frequency characteristics of various degradations but also effectively discriminates between different degradations, we perform t-SNE visualization on standard “noise-rain-haze” setting on the latent degradation representations obtained by PromptIR [17], Uformer [43] (i.e., the Rformer without frequency modulation), and our method. The results are shown in Fig. 9. Without frequency modulation, Uformer [43] fails to provide a discriminative capability superior to PromptIR[17]. However, with the incorporation of frequency modulation, our method effectively distinguish between different degradation types.
IV-F Ablation Studies
In this section, we conduct all ablation experiments in Table V on the standard “noise-rain-haze” setting to validate the components of the proposed network architecture.
Intra- & Inter-Band attention. We conduct this ablation experiment to demonstrate the effectiveness of the proposed Intra-Band & Inter-Band attention mechanism and the rationale behind the separation strategy of frequency bands. We replace the proposed attention mechanism with the origin window-based multi-head self-attention mechanism proposed by Swin Transformer [48]. As a result, a significant performance drop of 0.48 dB PSNR in dehazing task can be found in ”w.o. intra-inter” in Table V, as the origin attention mechanism struggles to capture the interactions among different frequency bands.
Frequency Decomposition. To illustrate the importance of frequency decomposition in the head layers of the Encoder to suppress the inherent bias of Transformer-like model towards different frequency information, we replaced the frequency decomposition layers in the Encoder with a simple MLP layer to achieve the same goal of dividing the features into $L$ sub-maps. In the ”w.o. Decompose” column of Table V, we observed a certain degree of performance degradation across all 3 degradations. Specifically, deraining and dehazing experienced a decrease of 2 dB and 0.82 dB, respectively, in terms of PSNR. The significant performance drop indicates that without frequency decomposition, there is a pronounced negative interactions between different frequency bands, thereby impairing the performance of the Decoder when embedding degradation representations.
TABLE V: Ablation staudy. ”w.o. intra-inter” means replacing the dual MSA with the regular MSA;”w.o. Decompose” means removing the initial frequency decomposition in the encoder and replacing the dual MSA with the regular MSA..
Frequency-aware Denoise Derain Dehaze
Decompose Intra- & Inter-Band BSD68 $(\sigma=15)$ BSD68 $(\sigma=25)$ BSD68 $(\sigma=50)$ Rain100L SOTS
34.59/0.942 31.84/0.900 28.48/0.814 37.42/0.980 28.17/0.968
$\checkmark$ 34.59/0.941 31.83/0.900 28.48/0.813 37.45/0.980 28.72/0.971
$\checkmark$ 34.44/0.940 31.71/0.898 28.35/0.809 35.50/0.970 28.38/0.968
$\checkmark$ $\checkmark$ 34.59/0.941 31.83/0.900 28.46/0.814 37.50/0.980 29.20/0.972
TABLE VI: Comparison to the state-of-the-art on 5 tasks All-in-one setting. Our method consistently achieves state-of-the-art results.
Method Params. Dehaze Derain Denoise Deblur Low-light Average
SOTS Rain100L BSD68 $(\sigma=25)$ GOPro LOL
DL[77] 2M 20.54/0.826 21.96/0.762 23.09/0.745 19.86/0.672 19.83/0.712 21.05/0.743
Transweather[79] 38M 21.32/0.885 29.43/0.905 29.00/0.841 25.12/0.757 21.21/0.792 25.22/0.836
TAPE[80] 1M 22.16/0.861 29.67/0.904 30.18/0.855 24.47/0.763 18.97/0.621 25.09/0.801
AirNet[15] 9M 21.04/0.884 32.98/0.951 30.91/0.882 24.35/0.781 18.18/0.735 25.49/0.846
IDR[81] 15M 25.24/0.943 35.63/0.965 31.60/0.887 27.87/0.846 21.34/0.826 28.34/0.893
PromptIR[17] 36M 26.54/0.949 36.37/0.970 31.47/0.886 28.71/0.881 22.68/0.832 29.15/0.904
Gridformer[82] 34M 26.79/0.951 36.61/0.971 31.45/0.885 29.22/0.884 22.59/0.831 29.33/0.904
Ours 32M 27.04/0.961 37.51/0.978 31.77/0.893 29.13/0.884 23.02/0.840 29.69/0.911
TABLE VII: Comparison on real-world denoising. Our method achieves state-of-the-art performance, validating its effectiveness in handling complex real-world scenarios.
Method MIRNet[83] DeamNet[84] MPRNet[52] DAGL[85] Uformer[43] Restormer[42] SSAMAN[86] NAFNet[87] Ours
SIDD 39.72/0.959 39.47/0.957 39.71/0.958 38.94/0.953 39.77/0.959 40.02/0.960 40.08/0.962 40.30/0.961 40.38/0.971
IV-G Generalization analysis
Performance on 5 tasks All-in-one setting. Informed by recent studies, including Gridformer [82] and InstructIR [19], we extended the existing four degradation types by incorporating the LOL dataset [74] for low-light degradation to further validate the efficacy of the proposed method. A comprehensive experiment encompassing all five degradation types was then conducted. Furthermore, we compared our method against a broader range of literature focusing on the same task setting. The experimental results are detailed in Table VI. As demonstrated, our method achieves a performance gain of 0.36 dB over the current state-of-the-art method.
Performance on real-world denoising. To further demonstrate that our method is not limited to synthetically injected noise, we followed the experimental setup of Restormer[42] and conducted experiments on the real-world image denoising dataset SIDD [88]. Specifically, without leveraging any external data, we trained and tested our model directly on the SIDD dataset. As shown in Table VII, our method achieves state-of-the-art performance, validating its effectiveness in handling complex real-world scenarios.